If P.E of bohr's hydrogen atom in its ground state is zero. Then its total energy in the first excited sate is?
 
Answer provided Is 23.8eV

Grade:12

1 Answers

Umakant biswal
7 years ago
dear yash 
i can help you with the correct approach to this question , rest u need to apply the formula and find out . Total Energy of electron, Etotal = Potential energy (PE) + Kinetic energy (KE) 
For an electron revolving in a circular orbit of radius, r around a nucleus with Z positive charge, 
PE = -Ze^2/r
KE = Ze^2/2r
hence 
Etotal = (-Ze^2/r) + (Ze^2/2r) = -Ze^2/2r
And for H atom, Z = 1
therefore , 
Etotal = -e^2/2r
first excited state is 2nd orbit. 
HOPE IT CLEARS NOW 
​all the best 
regards